# No-quenching baseline for energy loss signals in oxygen-oxygen collisions

### v1

@@ 3.3 Scale, shower, and hadronization uncertainties
The factorization and renormalization scales are not physical and the dependence of the cross section on them is the consequence of the truncated perturbative series. Therefore, varying $\mu_{R,F}$ is commonly used to estimate the uncertainties due to missing higher-order terms [^5] Note that scale variations do not have a probabilistic interpretation of uncertainties. Instead, Bayesian models can be used to estimate the confidence intervals of missing higher-order terms Duhr:2021mfd .. We choose them to be equal to the sum of outgoing parton transverse momenta $\mu_{F,R}=\frac{1}{2}\sum_{i}p_{T,i}$ and follow a 9-point variation technique to estimate the scale uncertainties by varying each of them by a factor of two up and down. In particular, we vary the scales correlatedly for all cross sections in $I_{\mathrm{AA}}$. Figure 8 shows the scale uncertainties of the away-side $I_{\rm AA}$ for LHC kinematics. They are around a few percent across all $p_{T}$, smaller than most nPDF uncertainty. In the low momentum region, nPDF and scale uncertainties are similar. Near-side hadrons, lower center-of-mass energies, and different $p_{T,\min}^{j}$ trigger cuts (not shown) have similar scale uncertainties.
@@
The trigger jet cross section is infrared and collinear safe (IRC), and robust with respect to the hadronization model. However, the distribution of charged hadrons can be sensitive to the details of the partonic shower and the non-perturbative hadronization effects. Here, we estimate the shower and hadronization uncertainties by comparing computations with two LO event generators: Pythia8.312 and Herwig7.3.0, which use different shower and hadronization models. The left panel of Fig. 9 shows $Y_{\rm AA}$ computed for partons, charged and all hadrons. Trigger jets were reconstructed on all hadrons or partons. The parton number is not an IRC-safe quantity. Partons fragment into lower momentum hadrons; therefore, partonic $Y_{pp}$ is shifted to larger momentum compared with hadrons. The all-hadron $Y_{pp}$ differs between Pythia and Herwig, roughly by a constant factor. Compared with the spectra including only charged hadrons, the all-hadron spectra are larger by a factor of $1.2$ $1.3$. The trends are similar for away- and near-side particles (not shown).
Finally, the right panel of Fig. 9 shows $I_{\rm AA}$ with drastically reduced hadronization differences. We observed that there is no difference between charged and all-hadron $I_{\rm AA}$. However, there is a systematic difference between Pythia and Herwig results, the latter being up to $5\%$ suppressed. In the low momentum region, the difference is similar in size to that of nPDF uncertainties. We note the positive slope in this region for $I_{\rm AA}$ in LO computations, which is absent in NLO computations, indicating that further studies of shower and hadronization uncertainties might be needed. Hadronization uncertainties can be reduced by measuring recoiling jets instead of hadrons, i.e., dijets ATLAS:2023zfx ; ALICE:2023plt ; ATLAS:2024jtu ; Li:2024uzk .
Figure 9: Left: Jet triggered semi-inclusive particle spectrum $Y_{\rm OO}$ normalised to Pythia charged. Right: The jet triggered semi-inclusive nuclear modification factor $I_{\rm AA}$ for partons, charged and all hadrons. Bands indicate statistical uncertainties.
@@ 4 Hadron-triggered jet nuclear modification factor $I_{\rm AA}^{j}$
No-quenching predictions of $I_{\mathrm{OO}}^{j}$ with $p_{T,\mathrm{min}}^{h}=12\,\text{GeV}$ at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$ are presented in the left panel of Fig. 10. The EPPS21 nPDF uncertainties are growing with momentum from $5\%$ up to $15\%$. The central values show significant deviations from unity ranging from $5\%$ to approximately $10\%$. Even within large nPDF uncertainties, the EPPS21 baseline shows deviations from unity. At lower $p_{T}^{j}$ TUJU21 nPDF uncertainties are around one percent and become roughly $5\%$ at larger momentum. Although central values computed with TUJU21 are closer to unity, deviations are again visible.
A correlation analysis of the trigger cross section $\sigma^{h}$ and the coincidence cross section $\sigma^{j+h}$ in the form of the Pearson correlation coefficient is presented in the right panel of Fig. 10. While an almost perfect correlation is apparent in the lowest $p_{T}^{j}$ bin, it declines quickly for higher momenta. Cross sections obtained with TUJU21 even become anti-correlated for $p_{T}^{j}>100\,\text{GeV}$. The loss of correlation accompanies growing nPDF uncertainties for $I^{j}_{\mathrm{OO}}$. In comparison to jet-triggered results, correlations are lost more quickly. This can be traced back to the absence of a $p_{T}^{j}$ window in which the two cross sections necessarily probe similar Bjorken-$x$. The plotted transverse momentum range of the jet is larger than the trigger threshold $p_{T,\mathrm{min}}^{h}$. However, the trigger hadrons are dominated by the fragmentation of near-threshold momentum jets. Therefore, the correlation is quickly lost for jets with momentum that is twice the threshold.
@@
Results at $\sqrt{s_{\mathrm{NN}}}=200\,\mathrm{GeV}$, shown in Fig. 11, agree with earlier observations. Again, we find that at lower collision energies, nPDF uncertainties using TUJU21 are larger than for EPPS21. Overall, errors for hadron-triggered $I_{\mathrm{AA}}$ tend to be larger than for the jet-triggered $I_{\mathrm{AA}}$. We see that, in this case, there is a significant discrepancy between the baselines with EPPS21 and TUJU21 nPDFs, which emphasizes the need to consider multiple nPDF parametrizations for reliable baseline prediction.
@@ 5 Conclusions and outlook
We find that the uncertainty in oxygen nPDFs leads to up to $10\%$ uncertainties on jet nuclear modification factor $R_{\mathrm{AA}}^{j}$ at LHC collision energies. The central value is consistent with unity within nPDF uncertainties in the studied momentum range. Scale, hadronization, and shower uncertainties are negligible compared with those from nPDFs. At RHIC collision energies, $R_{\mathrm{AA}}^{j}$ uncertainties are similar in size but with clear suppression at large jet momentum due to the nPDFs. Different nPDF extractions lead to different central values, but mostly consistent nPDF uncertainty bands. These uncertainties and deviations from $R_{\rm AA}=1$ must be taken seriously when interpreting the magnitude of energy loss in measurements.
@@
The nPDF uncertainties can be drastically reduced by considering a double ratio of cross sections, e.g., jet-triggered semi-inclusive $I_{\rm AA}$. However, we find such uncertainty cancelation (down to percent) only for recoiling hadrons below the jet trigger momentum. At higher momentum, the $I_{\rm AA}$ uncertainties increase substantially because the trigger jets and recoiling high momentum hadrons probe nPDFs at different Bjorken-$x$. Although, at LHC, $I_{\rm AA}$ remains consistent with unity within uncertainty bands, central values of different nPDF sets can show opposing trends. This underlies the importance of systematically comparing several nPDFs for robust baseline computation. At RHIC energies, we observe a significant suppression (up to $20\%$) of $I_{\rm AA}$ for large hadron momentum due to nPDFs. We also show that scale, parton shower, and hadronization uncertainties contribute additional percent level uncertainty in the low momentum region.
@@
For hadron-triggered semi-inclusive jet $I_{\rm AA}$, the nPDF uncertainty is observed to cancel close to the hadron trigger momentum, but it grows rapidly at larger momenta. We note that at LHC energies hadron-triggered $I_{\rm AA}$ is above unity, even considering the nPDF uncertainties. Therefore unity clearly cannot be used as a baseline to quantify the potential energy loss signal (see also Ref. He:2024rcv ). We note that at RHIC energies, predictions between different nPDF sets do not agree, which prevents us from making precise baseline predictions for these kinematics.
@@
In conclusion, we presented state-of-the-art predictions for the no-quenching baseline for jet $R_{\mathrm{AA}}^{j}$ as well as jet-, and hadron-triggered $I_{\rm AA}$ observables with robust estimation of theoretical uncertainties in oxygen-oxygen collisions [^6] We expect near identical results in other light ion systems, e.g., neon-neon collisions.. Whether an experimental discovery of energy loss in such a small system can be done will also crucially depend on the experimental precision of these measurements. With the expected integrated luminosity of $0.5\,\text{nb}^{-1}$ Brewer:2021kiv , the statistics at LHC could be sufficient for a precise jet-triggered hadron $I^{h}_{\rm AA}$ measurement, see Appendix C. Although minimum bias OO collisions do not have centrality selection uncertainty, $R_{\mathrm{AA}}$ and $I_{\mathrm{AA}}$-type measurements require a suitable $pp$ reference at the same collision energy. If a reference measurement at $\sqrt{s_{\text{NN}}}=6.8\,\text{TeV}$ is not available, alternative strategies of reference spectra interpolation or ratio of spectra at different energies could be considered Brewer:2021tyv . However, the estimation of experimental and theoretical uncertainties of such approaches is beyond the scope of the current paper. Finally, a discovery is only possible if a sufficiently large signal is present. Although some predictions for $R_{\mathrm{AA}}$-type observables have been computed in the past Huss:2020dwe ; Zakharov:2021uza ; Ke:2022gkq , a systematic comparison of the sensitivity of different observables to medium modifications would be very valuable in guiding experimental and theoretical developments. We leave such computations for future work.
@@ Acknowledgements.
The authors thank Valentin Hirschi, Alexander Huss, Peter Jacobs, and Petja Paakkinen for useful discussions. This work is supported by the DFG through Emmy Noether Programme (project number 496831614) and CRC 1225 ISOQUANT (project number 27381115). We acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 39/963-1 FUGG (bwForCluster NEMO).
@@ Appendix A Event generator settings
Appendix B Absolute inclusive and semi-inclusive spectra
@@
After demonstrating that the setup is well capable of reproducing inclusive observables, Fig. 13 shows its performance for jet-triggered $Y_{\mathrm{AA}}^{h}$ using $p_{T,\mathrm{min}}=30\,\mathrm{GeV}$, defined in Eq. 4 and Eq. 6. For $p_{T}<2\,\mathrm{GeV}$, the numerical predictions agree well with the experiment. Above $2\,\mathrm{GeV}$, the computation overpredicts the number of produced hadrons. The overproduction is less severe at LO than at NLO. Consequently, the LO predictions are closer to experimental observation while still not in agreement. When computing the ratio $Y_{\mathrm{AA}}/Y_{pp}=I_{\mathrm{AA}}$, the difference in magnitude cancels. Near-side and away-side comparisons (left and right columns) show identical characteristics.
@@
Appendix C Projection of statistical uncertainties
@@
Figure 14: Projection of relative statistical uncertainties for trigger and coincidence cross sections at $\sqrt{s_{\mathrm{NN}}}=6.8\,\mathrm{TeV}$. Left: Jet-triggered measurement with trigger momentum cutoff $p_{T,\mathrm{min}}^{j}=30\,\mathrm{GeV}$ (upper panel) and $60\,\mathrm{GeV}$ (lower panel). Results are shown for trigger cross section $\sigma^{j}$ and away-side $\mathrm{d}\sigma^{h+j}/\mathrm{d}p_{T}$. Right: Hadron-triggered measurement with $p_{T,\mathrm{min}}^{h}=12\,\mathrm{GeV}$. Results are shown for trigger cross section $\sigma^{h}$ and $\mathrm{d}\sigma^{j+h}/\mathrm{d}p_{T}^{j}$.
@@
Figure 15: Projection of relative statistical uncertainties for trigger and coincidence cross sections at $\sqrt{s_{\mathrm{NN}}}=200\,\mathrm{GeV}$. Left: Jet-triggered measurement with trigger momentum cutoff $p_{T,\mathrm{min}}^{j}=15\,\mathrm{GeV}$. Results are shown for trigger cross section $\sigma^{j}$ and away-side $\mathrm{d}\sigma^{h+j}/\mathrm{d}p_{T}$. Right: Hadron-triggered measurement with $p_{T,\mathrm{min}}^{h}=12\,\mathrm{GeV}$. Results are shown for trigger cross section $\sigma^{h}$ and $\mathrm{d}\sigma^{j+h}/\mathrm{d}p_{T}^{j}$.

### v3

@@ 3.3 Scale, shower, and hadronization uncertainties
The factorization and renormalization scales are not physical and the dependence of the cross section on them is the consequence of the truncated perturbative series. Therefore, varying $\mu_{R,F}$ is commonly used to estimate the uncertainties due to missing higher-order terms [^3] Note that scale variations do not have a probabilistic interpretation of uncertainties. Instead, Bayesian models can be used to estimate the confidence intervals of missing higher-order terms Duhr:2021mfd .. We choose them to be equal to the sum of outgoing parton transverse momenta $\mu_{R,F}=\frac{1}{2}\sum_{i}p_{T,i}$ and follow a 9-point variation technique to estimate the scale uncertainties by varying each of them by a factor of two up and down. In particular, we vary the scales correlatedly for all cross sections in $I_{\mathrm{AA}}$. Figure 8 shows the scale uncertainties of the away-side $I_{\rm OO}$ for LHC kinematics [^4] In this plot the central value of nNNPDF3.0 corresponds to the $0^{\mathrm{th}}$ member of the nPDF set which is the average over all replicas.. They are around a few percent across all $p_{T}$, smaller than most nPDF uncertainty. In the low momentum region, nPDF and scale uncertainties are similar. Near-side hadrons, lower center-of-mass energies, and different $p_{T,\min}^{j}$ trigger cuts (not shown) have similar scale uncertainties.
@@
The trigger jet cross section is infrared and collinear safe (IRC), and robust with respect to the hadronization model. However, the distribution of charged hadrons can be sensitive to the details of the partonic shower and the non-perturbative hadronization effects. Since these are modeled phenomenologically, instead of a systematic uncertainty study, we estimate shower and hadronization dependence by comparing computations with two LO event generators: Pythia8.312 and Herwig7.3.0, which use different shower and hadronization models. The left panel of Fig. 9 shows $Y_{\rm OO}$ computed for partons, charged and all hadrons. Trigger jets were reconstructed on all hadrons or partons. The parton number is not an IRC-safe quantity. Partons fragment into lower momentum hadrons; therefore, partonic $Y_{\rm OO}$ is shifted to larger momentum compared with hadrons. The all-hadron $Y_{\rm OO}$ differs between Pythia and Herwig, roughly by a constant factor. Compared with the spectra including only charged hadrons, the all-hadron spectra are larger by a factor of $1.2$ $1.3$. The trends are similar for away- and near-side particles (not shown).
Finally, the right panel of Fig. 9 shows $I_{\rm OO}$ with drastically reduced hadronization differences. We observed that there is no difference between charged and all-hadron $I_{\rm OO}$. However, there is a systematic difference between Pythia and Herwig results, the latter being up to $5\%$ suppressed. In the low momentum region, the difference is similar in size to that of nPDF uncertainties. We note the positive slope in this region for $I_{\rm OO}$ in LO computations, which is absent in NLO computations, indicating that further studies of shower and hadronization uncertainties might be needed. Sensitivity to hadronization and shower models can be reduced by measuring recoiling jets instead of hadrons, i.e., dijets ATLAS:2023zfx ; ALICE:2023plt ; ATLAS:2024jtu ; Li:2024uzk .
Figure 9: Left: Jet triggered semi-inclusive particle spectrum $Y_{\rm OO}$ normalised to Pythia charged. Right: The jet triggered semi-inclusive nuclear modification factor $I_{\rm OO}$ for partons, charged and all hadrons. Bands indicate statistical uncertainties.
@@ 4 Hadron-triggered jet nuclear modification factor $I_{\rm AA}^{j}$
No-quenching predictions of $I_{\mathrm{OO}}^{j}$ with $p_{T,\mathrm{min}}^{h}=12\,\text{GeV}$ at $\sqrt{s_{\mathrm{NN}}}=6.8\,\text{TeV}$ are presented in the left panel of Fig. 10. The EPPS21 and similarly nNNPDF3.0 nPDF uncertainties are growing with momentum from $5\%$ up to $15\%$. The central values show significant deviations from unity ranging from $5\%$ to approximately $10\%$. Even within large nPDF uncertainties, the EPPS21 baseline shows deviations from unity. At lower $p_{T}^{j}$ TUJU21 nPDF uncertainties are around one percent and become roughly $5\%$ at larger momentum. Although central values of TUJU21 are closer to unity, deviations are again visible.
A correlation analysis of the trigger cross section $\sigma^{h}$ and the coincidence cross section $\sigma^{j+h}$ in the form of the Pearson correlation coefficient is presented in the right panel of Fig. 10. While an almost perfect correlation is apparent in the lowest $p_{T}^{j}$ bin, it declines quickly for higher momenta. Cross sections obtained with TUJU21 and nNNPDF3.0 even become anti-correlated for $p_{T}^{j}>75\,\text{GeV}$. The loss of correlation accompanies growing nPDF uncertainties for $I^{j}_{\mathrm{OO}}$. In comparison to jet-triggered results, correlations are lost more quickly. This can be traced back to that the trigger hadron originates from a jet with slightly higher $p_{T}^{j}$, not far from the threshold, while jets on the other side with much higher momentum will probe significantly different Bjorken-$x$. Therefore, the correlation is quickly lost for jets with momentum that is twice the threshold.
@@
Results at $\sqrt{s_{\mathrm{NN}}}=200\,\mathrm{GeV}$, shown in Fig. 11, agree with earlier observations. Again, we find that at lower collision energies, nPDF uncertainties using TUJU21 and nNNPDF3.0 are larger than for EPPS21. Overall, errors for hadron-triggered $I_{\mathrm{OO}}$ tend to be larger than for the jet-triggered $I_{\mathrm{OO}}$. We see that, in this case, there is a significant discrepancy between the baselines with EPPS21 and TUJU21 nPDFs, which emphasizes the need to consider multiple nPDF parametrizations for reliable baseline prediction.
@@ 5 Conclusions and outlook
We find that the uncertainty in oxygen nPDFs leads to up to $10\%$ uncertainties on jet nuclear modification factor $R_{\mathrm{OO}}^{j}$ at LHC collision energies. The central value is consistent with unity within nPDF uncertainties in the studied momentum range. Scale, hadronization, and shower uncertainties are negligible compared with those from nPDFs. At RHIC collision energies, $R_{\mathrm{OO}}^{j}$ uncertainties are similar in size but with clear suppression at large jet momentum due to the nPDFs. Different nPDF extractions lead to different central values, but mostly consistent nPDF uncertainty bands. These uncertainties and deviations from $R_{\rm OO}=1$ must be taken seriously when interpreting the magnitude of energy loss in measurements.
@@
The nPDF uncertainties can be drastically reduced by considering a double ratio of cross sections, e.g., jet-triggered semi-inclusive $I_{\rm AA}$. However, we find such uncertainty cancellation only for recoiling hadrons below the jet trigger momentum. At higher momentum, the $I_{\rm OO}$ uncertainties increase substantially because the trigger jets and recoiling high momentum hadrons probe nPDFs at different Bjorken-$x$. Although, at LHC, $I_{\rm OO}$ remains consistent with unity within uncertainty bands, central values of different nPDF sets can show opposing trends. This underlies the importance of systematically comparing several nPDFs for robust baseline computation. At RHIC energies, we observe a significant suppression (up to $20\%$) of $I_{\rm OO}$ for large hadron momentum due to nPDFs. We also show that scale uncertainties cancel significantly and predictions show little sensitivity to parton shower and hadronization models in the double ratio.
@@
For hadron-triggered semi-inclusive jet $I_{\rm OO}$, the nPDF uncertainty is observed to cancel close to the hadron trigger momentum, but it grows rapidly at larger momenta. We note that at LHC energies hadron-triggered $I_{\rm OO}$ is above unity, even considering the nPDF uncertainties. Therefore unity clearly cannot be used as a baseline to quantify the potential energy loss signal (see also Ref. He:2024rcv ). We note that at RHIC energies, predictions between different nPDF sets do not agree, which prevents us from making precise baseline predictions for these kinematics.
@@
In conclusion, we presented state-of-the-art predictions for the no-quenching baseline for jet $R_{\mathrm{AA}}^{j}$ as well as jet-, and hadron-triggered $I_{\rm AA}$ observables with robust estimation of theoretical uncertainties along with model dependencies in oxygen-oxygen collisions [^5] We expect near identical results in other light ion systems, e.g., neon-neon collisions.. Whether an experimental discovery of energy loss in such a small system can be done will also crucially depend on the experimental precision of these measurements. With the expected integrated luminosity of $0.5\,\text{nb}^{-1}$ Brewer:2021kiv , the statistics at LHC could be sufficient for a precise jet-triggered hadron $I^{h}_{\rm AA}$ measurement, see Appendix D. Although minimum bias OO collisions do not have centrality selection uncertainty, $R_{\mathrm{AA}}$ and $I_{\mathrm{AA}}$-type measurements require a suitable $pp$ reference at the same collision energy. If a reference measurement at $\sqrt{s_{\text{NN}}}=6.8\,\text{TeV}$ is not available, alternative strategies of reference spectra interpolation or ratio of spectra at different energies could be considered Brewer:2021tyv . However, the estimation of experimental and theoretical uncertainties of such approaches is beyond the scope of the current paper. Finally, a discovery is only possible if a sufficiently large signal is present. Although some predictions for $R_{\mathrm{AA}}$-type observables have been computed in the past Huss:2020dwe ; Zakharov:2021uza ; Ke:2022gkq , a systematic comparison of the sensitivity of different observables to medium modifications would be very valuable in guiding experimental and theoretical developments. We leave such computations for future work.
@@ Acknowledgements.
The authors thank Valentin Hirschi, Alexander Huss, Peter Jacobs, Petja Paakkinen, and Juan Rojo, for useful discussions. This work is supported by the DFG through Emmy Noether Programme (project number 496831614) and CRC 1225 ISOQUANT (project number 27381115). We acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant no INST 39/963-1 FUGG (bwForCluster NEMO).
Code and Data Availability. This article has associated code and data available on https://zenodo.org/records/15052297.
@@ Appendix A Event generator settings
Appendix B Comparing computations at NLO and LO
In some cases uncertainty computations at NLO with MadGraph_aMC@NLO can exhibit irregular behavior, leading to significant uncertainty increase. This phenomenon is illustrated in the left panel of Fig. 12, where we compare results for $R^{j}_{\mathrm{OO}}$ calculated using the NLO nNNPDF3.0 with both NLO and LO matrix elements. In particular, in certain $p_{T}$ bins both nPDF and scale uncertainties are significantly larger when using NLO matrix elements. Generally, the effect of going to NLO on central values and nPDF uncertainties are minor for the inclusive jet nuclear modification factor in Fig. 12 as well as semi-inclusive observables in Fig. 13 where we representatively show results for $I^{h}_{\mathrm{OO}}$. Since we did not manage to fully resolve the issue of irregular uncertainty behavior, and we consider it having numerical origin, when using nNNPDF3.0 in the main text, we show results with LO matrix elements.
Figure 12: Comparing NLO with LO calculations of inclusive jet $R^{j}_{\mathrm{OO}}$ using nNNPDF3.0 (left) and EPPS21 (right) nPDFs at $\sqrt{s_{\mathrm{NN}}}=6.8\,\mathrm{TeV}$.
Figure 13: Comparing NLO with LO calculations of semi-inclusive jet-triggered $I^{h}_{\mathrm{OO}}$ using nNNPDF3.0 (left) and EPPS21 (right) nPDFs at $\sqrt{s_{\mathrm{NN}}}=6.8\,\mathrm{TeV}$.
Appendix C Absolute inclusive and semi-inclusive spectra
@@
After demonstrating that the setup is well capable of reproducing inclusive observables, Fig. 15 shows its performance for jet-triggered $Y_{\mathrm{OO}}^{h}$ using $p_{T,\mathrm{min}}=30\,\mathrm{GeV}$, defined in Eq. 4 and Eq. 6. For $p_{T}<2\,\mathrm{GeV}$, the numerical predictions agree well with the experiment. Above $2\,\mathrm{GeV}$, the computation overpredicts the number of produced hadrons. The overproduction is less severe at LO than at NLO. Consequently, the LO predictions are closer to experimental observation while still not in agreement. When computing the ratio $Y_{\mathrm{AA}}/Y_{pp}=I_{\mathrm{AA}}$, the difference in magnitude cancels. Near-side and away-side comparisons (left and right columns) show identical characteristics.
@@
Appendix D Projection of statistical uncertainties
@@
Figure 16: Projection of relative statistical uncertainties for trigger and coincidence cross sections at $\sqrt{s_{\mathrm{NN}}}=6.8\,\mathrm{TeV}$ using cross sections obtained with EPPS21. Left: Jet-triggered measurement with trigger momentum cutoff $p_{T,\mathrm{min}}^{j}=30\,\mathrm{GeV}$ (upper panel) and $60\,\mathrm{GeV}$ (lower panel). Results are shown for trigger cross section $\sigma^{j}$ and away-side $\mathrm{d}\sigma^{h+j}/\mathrm{d}p_{T}$. Right: Hadron-triggered measurement with $p_{T,\mathrm{min}}^{h}=12\,\mathrm{GeV}$. Results are shown for trigger cross section $\sigma^{h}$ and $\mathrm{d}\sigma^{j+h}/\mathrm{d}p_{T}^{j}$.
@@
Figure 17: Projection of relative statistical uncertainties for trigger and coincidence cross sections at $\sqrt{s_{\mathrm{NN}}}=200\,\mathrm{GeV}$ using cross sections obtained with EPPS21. Left: Jet-triggered measurement with trigger momentum cutoff $p_{T,\mathrm{min}}^{j}=15\,\mathrm{GeV}$. Results are shown for trigger cross section $\sigma^{j}$ and away-side $\mathrm{d}\sigma^{h+j}/\mathrm{d}p_{T}$. Right: Hadron-triggered measurement with $p_{T,\mathrm{min}}^{h}=12\,\mathrm{GeV}$. Results are shown for trigger cross section $\sigma^{h}$ and $\mathrm{d}\sigma^{j+h}/\mathrm{d}p_{T}^{j}$.